# LongRAG: Enhancing Retrieval-Augmented Generation with Long-context LLMs

### v2

@@ Abstract
In traditional RAG framework, the basic retrieval units are normally short. The common retrievers like DPR normally work with 100-word Wikipedia paragraphs. Such a design forces the retriever to search over a large corpus to find the ‘needle’ unit. In contrast, the readers only need to extract answers from the short retrieved units. Such an imbalanced ‘heavy’ retriever and ‘light’ reader design can lead to sub-optimal performance. In order to alleviate the imbalance, we propose a new framework LongRAG, consisting of a ‘long retriever’ and a ‘long reader’. LongRAG processes the entire Wikipedia into 4K-token units, which is 30x longer than before. By increasing the unit size, we significantly reduce the total units from 22M to 600K. This significantly lowers the burden of retriever, which leads to a remarkable retrieval score: answer recall@1=71% on NQ (previously 52%) and answer recall@2=72% (previously 47%) on HotpotQA (full-wiki). Then we feed the top-k retrieved units ($\approx$ 30K tokens) to an existing long-context LLM to perform zero-shot answer extraction. Without requiring any training, LongRAG achieves an EM of 62.7% on NQ and 64.3% on HotpotQA (full-wiki), which is on par with the (fully-trained) SoTA model. Our study offers insights into the future roadmap for combining RAG with long-context LLMs.
LongRAG: Enhancing Retrieval-Augmented Generation
with Long-context LLMs
Ziyan Jiang University of Waterloo ziyanjiang528@gmail.com                        Xueguang Ma University of Waterloo x93ma@uwaterloo.ca                        Wenhu Chen University of Waterloo wenhuchen@uwaterloo.ca
@@ 1 Introduction
Figure 1: Traditional RAG vs. LongRAG. (Up) Traditional RAG operates on short retrieval units, where the retriever needs to scan over a massive amount of units to find the relevant piece. In contrast, LongRAG operates on long retrieval units (30x longer). Retriever has a much less workload, which significantly boosts the recall score. LongRAG fully exploits the ability of long-context language models to achieve strong performance.
@@
Retrieval-Augmented Generation (RAG) methods have long been employed to enhance large language models (LLMs) Mialon et al. (2023). Knowledge in the form of natural language can be entirely offloaded from the parametric knowledge of LLMs by leveraging a standalone retrieval component from an external corpus. The existing RAG framework tends to use short retrieval units, such as 100-word passages in popular open-domain question-answering tasks Chen et al. (2017); Lewis et al. (2020); Karpukhin et al. (2020). The retriever is tasked with finding the “needle” (i.e. the precise tiny retrieval unit) from the “haystack” (i.e. the massive corpus with tens of millions of information units). Subsequently, the retrieved units are passed to the reader to generate the final response. On the contrary, the reader only needs to extract answers from these retrievals, which is a fairly easy task. This kind of imbalanced design, with a “heavy” retriever and a “light” reader, puts too much pressure on the retriever. Therefore, existing RAG models Izacard and Grave (2020b) have to recall huge amounts of units, such as the top-100/200, combined with additional re-ranker to achieve the best performance. Moreover, short retrieval units can lead to semantic incompleteness due to document truncation. This can lead to information loss, ultimately hurting the end performance. This design choice was made in an era when NLP models were heavily restricted by their ability to handle long contexts. With the recent advances in long-context language models, the retriever and reader can potentially handle up to 128K or even millions of tokens as input Reid et al. (2024); Achiam et al. (2023).
In this paper, we propose to revisit this design choice for open-domain question answering and propose the LongRAG framework as a solution to balance the workload between the retriever and the reader, as illustrated in Figure 1. There are three important designs in our novel framework:
  1. 1.
@@
    Long Retrieval Unit: By using entire Wikipedia documents or grouping multiple related documents, we can construct long retrieval units with more than 4K tokens. This design could also significantly reduce the corpus size (number of retrieval units in the corpus). Then, the retriever’s task becomes much easier with more complete information.
  2. 2.
@@
    Long Retriever: The long retriever will identify coarse relevant information for the given query by searching through all the long retrieval units in the corpus. Only the top 4 to 8 retrieval units (without re-ranking) are used for the next step.
  3. 3.
@@
These three novel designs significantly boost the overall performance of RAG on open-domain question-answering tasks like NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). LongRAG has several advantages: 1) It does not require additional re-rankers and the best results can be attained by only considering the top 4-8 retrieved units. 2) The long retrieval unit amalgamates comprehensive information from related documents, which can be used directly to answer multi-hop questions without iterative retrieval.
@@
In our experiments, we adopt off-the-shelf retrievers like BGE Xiao et al. (2023) and readers like Gemini-1.5-Pro (Reid et al., 2024) or GPT-4o OpenAI (2024) without any tuning on NQ or HotpotQA. In our experiments, we reduce the NQ corpus size from 22M to 600K document units, which improves the answer recall@1 from 52% (DPR) to 71%. Similarly, we reduce the HotpotQA corpus size from 5M to 500K, which improves the recall@2 from 47% (DPR) to 72%. The improvement in retriever can significantly benefit the reader model. By exploiting the long-context understanding ability of GPT-4o, LongRAG can achieve an EM of 62% on NQ and 64% on HotpotQA. These results could be comparable to the strongest fully trained RAG models like Atlas Izacard et al. (2022) and MDR Xiong et al. (2020b).
@@
We perform ablation studies in subsection 4.4 to prove why longer retrieval units are necessary. Given a budget of 40K recall tokens, with ‘short retriever units’, we can increase the number of recalled units to reach a marvelously high recall score (91% for recall@200). However, the end performance dips significantly due to the huge amount of ‘hard negatives’, which confuses the reader. With ‘long retriever units’, we observe an entirely different trend. As we recall more units (from 1 to 8 units), both the recall and end performance will increase or plateau. The impact of ‘hard negative’ is much less severe in LongRAG. It shows that LongRAG can better exploit the advances in the long-context LLMs (reader). As the long-context methods evolve, the performance of LongRAG will continue to improve. Therefore, we believe the modern RAG systems should re-consider the granularity of their retrieval units to exploit the advantages of the current long-context LLMs.
Meanwhile, there is still room for improvement in our framework, particularly the need for stronger long embedding models, as shown in Table 3. Additionally, more general methods to formulate long retrieval units beyond hyperlinks will be helpful.
2 Related Work
2.1 Retrieval-Augmented Generation.
Augmenting language models with information retrieved from large corpora has become a popular and effective approach for knowledge-intensive tasks, particularly open-domain question answering. The predominant architecture follows a retriever-reader style Chen et al. (2017); Guu et al. (2020), where the input query retrieves information from a corpus, and a language model uses this information as additional context to make a final prediction. Recent work has focused on improving the retriever (Karpukhin et al., 2020; Xiong et al., 2020a; Qu et al., 2020; Xiong et al., 2020b; Khalifa et al., 2023), enhancing the reader (Izacard and Grave, 2020b; Cheng et al., 2021; Yu et al., 2021; Borgeaud et al., 2022), fine-tuning the retriever and reader jointly (Yu, 2022; Izacard et al., 2022; Singh et al., 2021; Izacard and Grave, 2020a), and integrating the retriever with the black-box language model (Yu et al., 2023; Shi et al., 2023; Trivedi et al., 2022). However, the impact of document granularity on the effectiveness and efficiency of the retrieval-augmented generation pipeline remains underexplored.
2.2 Long Context Large Language Models.
The effectiveness of Transformer-based models is hindered by the quadratic increase in computational cost relative to sequence length, especially when dealing with long context inputs. In order to solve this issue, different approaches have been proposed to mitigate computational issues, including sliding memory window and chunk segmentation (Hao et al., 2022; Ratner et al., 2023; Zhu et al., 2024b). FlashAttention Dao et al. (2022) has also been a pivotal strategy to significantly reduce the memory footprint to almost linear w.r.t sequence length.
To enable length extrapolation, RoPE Su et al. (2021) and AliBI Press et al. (2021) position encodings have shown potential to enable length extrapolation, which have been widely used in the literature. Recent endeavors have explored diverse strategies to tackle this challenge, which is mainly Position reorganization (Jin et al., 2024; An et al., 2024), Position interpolation (Chen et al., 2023a; Peng et al., 2023; Liu et al., 2024). Furthermore, alternative architectures beyond the Transformer have been explored to handle long inputs more naturally. These diverse approaches claim that they can enhance the capabilities of LLMs in processing long context inputs more efficiently.
2.3 Long Context Embedding
Recent efforts also increased the context length for embedding models, extending the supported text snippet length from a limit of 512 tokens to 32k tokens. Typically, the development of long-context embedding models involves first obtaining a long-context backbone model. This can be achieved either by pre-training with long inputs from scratch Günther et al. (2023); Nussbaum et al. (2024); Chen et al. (2024) or by utilizing existing large language models that support longer context Wang et al. (2023). Additionally, some works extend the capabilities of existing embedding models to handle long contexts by applying LLM content window extension methods on embedding models Zhu et al. (2024a); Peng and Quesnelle (2023), or by employing state-space encoder models Saad-Falcon et al. (2024).
@@ 3 LongRAG
Our proposed LongRAG framework is comprised of two components: the Long Retriever and the Long Reader. An illustrative example of these two components are depicted in Figure 2.
Figure 2: LongRAG example. On the left side, it shows that the long retrieval unit is grouped by Wikipedia documents through hyperlinks. Each retrieval unit contains an average of 4K tokens, corresponding to multiple related documents. On the right side, it shows a multi-hop question answer test case from HotpotQA. The final result can be achieved by using only a few retrieval units, which is then fed into a long reader.

### v3

@@ Abstract
In traditional RAG framework, the basic retrieval units are normally short. The common retrievers like DPR normally work with 100-word Wikipedia paragraphs. Such a design forces the retriever to search over a large corpus to find the “needle” unit. In contrast, the readers only need to generate answers from the short retrieved units. The imbalanced “heavy” retriever and “light” reader design can lead to sub-optimal performance. The loss of contextual information in the short, chunked units may increase the likelihood of introducing hard negatives during the retrieval stage. Additionally, the reader might not fully leverage the capabilities of recent advancements in LLMs. In order to alleviate the imbalance, we propose a new framework LongRAG, consisting of a “long retriever” and a “long reader”. In the two Wikipedia-based datasets, NQ and HotpotQA, where the average document size is less than 1K tokens, LongRAG processes the entire Wikipedia corpus into 4K-token units by grouping related documents, making these units 30 times longer than before. By increasing the unit size, we significantly reduce the total number of units from 22M to 600K. This greatly reduces the burden on the retriever, resulting in strong retrieval performance with only a few (less than 8) top units. Compared to traditional RAG, which may require hundreds of short units to achieve similar retrieval performance, our approach minimizes the likelihood of retrieving hard negatives while maintaining semantic integrity of each unit. Then we feed these retrieved units ($\approx$ 30K tokens) to an existing long-context LLM to perform zero-shot answer generation. Without requiring any training, LongRAG achieves an EM of 62.7% on NQ and 64.3% on HotpotQA, which are on par with the (fully-trained) SoTA model. Furthermore, we test on two non-Wikipedia-based datasets, Qasper and MultiFieldQA-en, where the average document length is already above 4K tokens. LongRAG processes each individual document as a single (long) unit rather than chunking them into smaller units. By doing so, we achieve an F1 score of 25.9% on Qasper (previously 22.5%) and 57.5% on MultiFieldQA-en (previously 51.2%). Our study offers insights into the future roadmap for combining RAG with long-context LLMs.
@@ 1 Introduction
Figure 1: Traditional RAG vs. LongRAG. (Up) Traditional RAG operates on short retrieval units, where the retriever needs to scan over a massive amount of units to find the relevant piece. In contrast, LongRAG operates on long retrieval units (30x longer). The retriever of LongRAG has a significantly reduced workload, achieving strong retrieval quality by leveraging only a few top units without the need for additional ranking mechanisms or other complex components. LongRAG could fully exploit the ability of long-context language models to achieve strong performance. (Down) QA performance compared with other methods on the NQ dataset and the HotpotQA dataset.
@@
Retrieval-Augmented Generation (RAG) methods have long been employed to enhance large language models (LLMs) (Mialon et al., 2023). Knowledge in the form of natural language can be entirely offloaded from the parametric knowledge of LLMs by leveraging a standalone retrieval component from an external corpus. The existing RAG framework tends to use short retrieval units, such as 100-word passages in popular open-domain question-answering tasks (Chen et al., 2017; Lewis et al., 2020; Karpukhin et al., 2020). The retriever is tasked with finding the “needle” (i.e. the precise tiny retrieval unit) from the “haystack” (i.e. the massive corpus with up to tens of millions of information units). Subsequently, the retrieved units are passed to the reader to generate the final response. On the contrary, the reader only needs to extract answers from these retrievals, which is a fairly easy task. This kind of imbalanced design, with a “heavy” retriever and a “light” reader, puts too much pressure on the retriever. Therefore, existing RAG models (Izacard & Grave, 2020b) have to recall huge amounts of units, such as the top-100/200, combined with additional re-ranker to achieve the best performance. Moreover, short retrieval units can lead to semantic incompleteness due to document truncation. This can result in the loss of contextual information, which may ultimately harm overall performance. This design choice was made in an era when the reader models were heavily restricted by their ability to handle long and contexts. With the recent advances in long-context language models, the reader can potentially handle up to 128K or even millions of tokens as input (Reid et al., 2024; Achiam et al., 2023). In this paper, we propose to revisit this design choice for open-domain question answering and propose the LongRAG framework as a solution to balance the workload between the retriever and the reader, as illustrated in Figure 1. There are three important designs in our novel framework:
@@
Long Retrieval Unit: By using entire documents or grouping multiple related documents, we can construct long retrieval units with more than 4K tokens. This design could also significantly reduce the corpus size (number of retrieval units in the corpus). This makes the retriever’s task much easier by providing more complete information, allowing the retriever’s architecture to be simplified without the need for additional re-rankers or iterative retrieval.
@@
Long Retriever: The long retriever will identify coarse relevant information for the given query by searching through all the long retrieval units in the corpus. Only a few top retrieval units (1 to 8 retrieval units in the four datasets we tested on), without re-ranking, are used for the next step. Compared to retrieving hundreds of short units, the long retriever only needs to retrieve a few candidates, which significantly reduces the likelihood to encounter hard negatives (it will confuse the reader).
@@
These three novel designs significantly boost the overall performance of RAG on open-domain question-answering tasks like NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), Qasper (Dasigi et al., 2021) and MultiFieldQA-en (Bai et al., 2023).
@@
In our experiments, we adopt off-the-shelf retrievers like BGE (Xiao et al., 2023) and readers like Gemini-1.5-Pro (Reid et al., 2024) or GPT-4o (OpenAI, 2024) without any further tuning. To demonstrate the generalizability of our proposed framework, we tested it on four datasets from different scenarios. First, we evaluate it on NQ and HotpotQA, which are Wikipedia-based dataset. The corpus of both datasets are composed of relatively short (averaging less than 1K tokens) but vast Wikipedia documents. By forming longer retrieval units through the grouping of multiple related documents, we reduce the NQ corpus size from 22M to 600K units, which improves the answer recall@1 from 52% (DPR) to 71%. Similarly, we reduce the HotpotQA corpus size from 5M to 500K, which improves the recall@2 from 47% (DPR) to 72%. By exploiting the long-context understanding ability of GPT-4o, LongRAG can achieve an EM of 62.7% on NQ and 64.3% on HotpotQA. These results could be comparable to the strongest fully trained RAG models like Atlas (Izacard et al., 2022) and MDR (Xiong et al., 2020b). Furthermore, we test on two non-Wikipedia-based datasets, Qasper and MultiFieldQA-en, where the corpus consists of relatively long documents averaging more than 4K tokens. LongRAG processes each entire document as a single unit rather than chunking them into smaller units. By doing so, we achieve an F1 score of 25.9% on Qasper (previously 22.5%) and 57.5% on MultiFieldQA-en (previously 51.2%).
@@
We perform ablation studies in subsection 3.5 to prove why longer retrieval units are necessary. Given a budget of 40K recall tokens, with “short retriever units”, we can increase the number of recalled units to reach a marvelously high recall score (91% for recall@200). However, the end performance dips significantly due to the huge amount of “hard negatives”, which confuses the reader. With “long retriever units”, we observe an entirely different trend. As we recall more units (from 1 to 8 units), both the recall and end performance will increase or plateau. The impact of “hard negative” is much less severe in LongRAG. It shows that LongRAG can better exploit the advances in the long-context LLMs (reader). As the long-context methods evolve, the performance of LongRAG will continue to improve. Therefore, we believe the modern RAG systems should re-consider the granularity of their retrieval units to exploit the advantages of the current long-context LLMs.
@@ 2 LongRAG
Our proposed LongRAG framework is comprised of two components: the Long Retriever and the Long Reader. Compared to traditional RAG, which operates on a large number of short retrieval units, LongRAG operates on long retrieval units, with only a few (typically fewer than 10) being fed into the reader. An illustrative example is shown in Figure 2.
Figure 2: LongRAG example. We form the long retrieval unit which is at least 4K tokens by using the entire document or grouping related documents, depending on the orinal docuemnt size. In this example, multiple Wikipedia documents are grouped through hyperlinks. This approach enables even multi-hop question-answering cases from HotpotQA to be addressed using only a few retrieval units, which are then fed into a long reader. Compared to traditional RAG, which retrieves hundreds of short units, our proposed LongRAG reduces the likelihood of retrieving hard negatives during the retrieval stage and more effectively leverages recent advances in long-context LLMs.